write an equation in slope-intercept form through (3,-1) with slope -3.
Answer by mananth(16946)

(3,-1) with slope -3.
plug the values to find the intercept b
y = mx +b
-1=-3*3+b
-1=-9+b
b=8
y=-3x+8